OpenStudy (anonymous):

Given f(x)=x-5/3, solve for f1(3) -1 14 18 21

OpenStudy (bibby):

does that say \(f^{-1}(3)\)?

OpenStudy (eliasperez191):

yes it does

OpenStudy (anonymous):

Yes

OpenStudy (bibby):

as always replace f(x) with y switch y and x solve for y our new y

OpenStudy (bibby):

\(y=x-\frac{5}{3}\longrightarrow x=y-\frac{5}{3}\\\)

OpenStudy (anonymous):

Nicely done, Bibby.

OpenStudy (bibby):

thanks buddy

OpenStudy (bibby):

oh shoot, I just realized you meant \(\frac{x-5}{3}\)

OpenStudy (anonymous):

Formatting is very important when asking for help.

OpenStudy (anonymous):

\[f(x)=\frac{ x-5 }{ 3 }\] This is what the equation so if I switch them it would look like: \[ Correct?

OpenStudy (bibby):

mhm \(y=\frac{x-5}{3}\longrightarrow x=\frac{y-5}{3}\)

OpenStudy (anonymous):

It messed up my equation -.-

OpenStudy (bibby):

looool

OpenStudy (bibby):

did you solve for the new y'?

OpenStudy (anonymous):

I'm working on solving it lol

OpenStudy (bibby):

oh, I thought you got mad because your messed up equation was the solved one btw you can use the draw button if you're feeling lazy

OpenStudy (anonymous):